# Near-field Beam Training with Sparse DFT Codebook

### v1

@@ Abstract
Extremely large-scale array (XL-array) has emerged as one promising technology to improve the spectral efficiency and spatial resolution of future sixth generation (6G) wireless systems. The upsurge in the antenna number antennas renders communication users more likely to be located in the near-field region, which requires a more accurate spherical (instead of planar) wavefront propagation modeling. This inevitably incurs unaffordable beam training overhead when performing a two-dimensional (2D) beam-search in both the angular and range domains. To address this issue, we first introduce a new sparse discrete Fourier transform (DFT) codebook, which exhibits the angular periodicity in the received beam pattern at the user, which motivates us to propose a three-phase beam training scheme. Specifically, in the first phase, we utilize the sparse DFT codebook for beam sweeping in an angular subspace and estimate candidate user angles according to the received beam pattern. Then, a central sub-array is activated to scan specific candidate angles for resolving the issue of angular ambiguity and identity the user angle. In the third phase, the polar-domain codebook is applied in the estimated angle to search the best effective range of the user. Finally, numerical results show that the proposed beam training scheme enabled by sparse DFT codebook achieves 98.67% reduction as compared with the exhaustive-search scheme, yet without compromising rate performance in the high signal-to-ratio (SNR) regime.
@@ I Introduction
Extremely large-scale array/surface (XL-array/surface) has been envisioned as one of the key ingredients to drive the evolution of the 6G wireless systems[1, 2, 3, 4]. Specifically, XL-arrays/surfaces with a significant number of antennas can be deployed at the base station (BS) to achieve ultra-high spectral efficiency and spatial resolution, hence accommodating meets the escalating demand for new applications such as spanning metaverse and digital twin[5, 6, 7]. The sharp increase in the number of antennas in high-frequency bands represents a qualitative paradigm shift in the electromagnetic (EM) propagation modeling, giving rise to the new near-field communications.
Particularly, different from far-field EM propagation which is simply approximated with planar waves, the near-field channel modeling necessitates the use of accurate spherical waves[8, 9, 10]. As such, near-field communications possess several unique properties in contrast to far-field communications. First, the spherical wavefront characteristic opens up the possibility of near-field beamfocusing, for which the beam energy can be concentrated at a specific location/region rather than a spatial angle typically for far-field beamforming[11, 12]. The beam-focusing capability of near-field communications enables XL-array to flexibly form highly directional beams in both the angle and range domain, and hence can be leveraged in various applications to improve the system performance, such as mitigating the inter-user interference, improving the accuracy of sensing and localization[13, 14, 15], and enhancing the charging efficiency of wireless power transfer[16]. Second, the rank of line-of-sight (LoS) channels for near-field multiple-input multiple-output (MIMO) communication systems can be larger than one, hence enhancing the spatial multiplexing gains [17]. In this paper, we propose an efficient near-field beam training scheme with an sparse DFT codebook by using the sparse antenna activation method, which significantly reduces the beam sweeping and thereby reduces the beam training overhead.
@@ I-A1 Near-Field Wireless Systems
Near-field communications bring new opportunities and challenges, which has motivated upsurging research interest. For example, in [11], the authors studied the beamfocusing design based on the fully-digital architectures, hybrid phase shifter-based precoders, and dynamic metasurface antenna architecture for XL-MIMO arrays. It is shown that near-field beamfocusing provides new DoF to mitigate the interference among users, even when they are located at the same spatial angle. In addition, a new concept of location division multiple access (LDMA) was proposed in [18], which exploits orthogonality of the near-field beamfocusing vectors in the range domain to serve different users from the same angle. The authors in [19] built a framework for analyzing and designing XL-MIMO systems with spatial non-stationarities. It was revealed that the performance of the proposed framework approaches that of the conventional whole array based designs but with lower complexity. In addition, a holographic metasurface antennas (HMAs) based multi-user system was investigated in [20], where the digital transmit precoder and the analog HMA weighting matrix were jointed optimized to minimize the transmit power. The authors in [21] considered a simultaneous wireless information and power transfer (SWIPT) system, where energy harvesting (EH) and information decoding receivers are located in the near- and far-field regions of the XL-array, respectively. The beam scheduling and power allocation were jointly optimized to maximize the weighted-sum power harvested at EH receivers[21]. Moreover, a directional modulation system was developed for near-field physical layer security[22]. Specifically, a fully analog precoding algorithm along with artificial noise and power allocation was proposed to realize secure transmission in both angular and range domains. Near-field sensing with XL-array was studied in [23], where the closed-form expressions of the Cramér-Rao Bounds for both angle and range estimations are derived.
@@ I-A2 Near-field Beam Training
In high-frequency bands, direct channel estimation methods may not be very effective due to severe path-loss and signal misalignment. As such, beam training is efficient in establishing initial links with high SNR for data transmission and channel state information (CSI) acquisition. However, due to the spherical wave propagation, near-field beam training is more challenging compared with far-field counterparts as it requires a joint beam search over both the angular and range domain. Particularly, conventional far-field beam training will experience significant performance loss in the near-field region due to the so-called energy-spread effect, for which the energy of a far-field beamformer is no longer steered towards one angle, but spread in multiple angles. Hence, the conventional far-field beam training method cannot be directly applied for near-field beam training. To address this issue, the work in [24] proposed a new codebook design in the polar domain for near-field communications, for which the angular domain is uniformly sampled whereas the range domain is non-uniformly sampled. One can simply invoke this codebook for exhaustive-search based near-field beam training. However, the overhead of the exhaustive-search based beam training scheme is the product of the number of antennas and range sampling points, which is prohibitively high for the implementation of XL-arrays. To reduce the overhead, the authors in [25] proposed a novel two-phase near-field beam training method. Leveraging a key observation that the true user angle approximately lies in the middle of an angular support with high received SNRs, this method first estimates the user angle by far-field DFT beams, and then estimates the user range with the polar-domain codebook in [24]. Besides, deep learning based near-field beam training methods have also been studied in [26] and [27] to reduce the overhead, where deep neural networks (DNN) are trained based on conventional far-field codebooks and near-field codebook, respectively. However, the training overhead of these methods scales linearly with the number of antennas, which is still unaffordable for communication systems. This issue motivates the design of hierarchical beam training schemes for near-field communications to reduce the training overhead to the logarithmic order, e.g., [28, 29, 30]. However, hierarchical beam training schemes suffer from inherent drawbacks. First, hierarchical methods require numerous feedback and neglect underlying transmission delays. Moreover, there exists error propagation problem due to the progressive beam training rounds [31].
@@ I-B Motivations and Contributions
The existing works on near-field beam training inevitably incur unacceptable overhead, while the off-grid channel estimation methods face highly computational complexity such as high dimensional matrix inversion. Moreover, the beam training overhead of these existing works scales linearly with the number of antennas apart from hierarchical beam training schemes, which suffers from several inherent drawbacks such as error propagation and user feed-back overhead [31]. Motivated by the above, this paper explores a new sparse DFT codebook to construct periodic received beam pattern at the user and then proposes an efficient three-phase beam training scheme, whose overhead scales with the square root of the number of antennas. The main contributions are summarized as follows.
  • First, we propose a novel sparse DFT codebook by sparsely activating the XL-array antennas and constructing the reduced DFT codebook with the equivalent sparse linear array (SLA). Specifically, the sparse DFT codebook consists of sparse far-field channel response vectors, which is sampled from far-field channel response vectors. Then, we characterize the received beam pattern at the near-field user when the sparse DFT codebook is used for beam sweeping. Interestingly, it is shown that the received beam pattern exhibits periodicity in the angular domain and there still exists the energy-spread effect. Then, we show that the user angle information is contained in a period of the received beam pattern at the user and can be estimated via a defined angular support.
  • Second, we propose a novel three-phase beam training scheme based on the sparse DFT codebook. Specifically, in the first phase, we utilize a small number of the sparse DFT codewords to sweep an angular subspace and estimate one candidate user angle according to the middle of the defined angular support. Then, in the second phase, we activate a central sub-array to resolve the angular ambiguity by virtue of the periodicity of the received beam pattern at the user. Finally, the polar-domain codebook is utilized to search the best user range in the estimated user angle.
  • Finally, extensive numerical results are presented to demonstrate the effectiveness of our proposed beam training scheme enabled by the sparse DFT codebook for near-field communications. It is shown that the proposed three-phase beam training scheme can achieve nearly the same performance with the exhaustive-search beam training method in the high-SNR regime, while reducing more than 98% of the training overhead. In the low-SNR regime, the proposed scheme suffers from slight performance loss, while the effective rate still significantly exceeds all benchmark schemes due to lower beam training overhead.
@@ I-C Organization and Notations
$D$ Array aperture size $M$ Number of antennas of the activated sub-array
@@
${{\mathcal{V}}}_{\rm{Sub}}$ DFT codebook for the activated sub-array $\bar{\mathcal{X}}_{\rm{Pol}}$ Polar-domain codebook
@@ II System Model
We consider a single-user XL-array downlink communication system, where the base station (BS) is equipped with a dense uniform linear array (ULA) with $N$ antennas. In this section, the near-field channel and signal model for the ULA are introduced.
@@ II-A Near-field Channel Model
which includes one LoS path $\mathbf{h}_{\rm LoS}^{H}$ and $L$ non-LoS (NLoS) paths. Herein, the parameters $r_{0}$ (${\bar{r}_{\ell}}$) and $\theta_{0}$ ($\bar{\theta}_{\ell}$) represent the range and spatial angle of the LoS ($\ell$-th NLoS) signal path. Moreover, the parameters $\beta$ and $\beta_{\ell}$ denote the LoS path and $\ell$-th NLoS path gain, respectively. Mathematically, $\beta$ can be modeled as $\beta=\sqrt{\frac{\kappa}{\kappa+1}}\frac{\sqrt{\beta_{0}}}{r_{{0}}}e^{-\frac{\jmath 2\pi r_{{0}}}{\lambda}}$, where $\kappa$ and $\beta_{0}$ represent the Rician factor and reference channel gain at a range of 1 m, respectively.
@@
with $r_{n}=\sqrt{r_{0}^{2}+{n}^{2}d_{0}^{2}-2r_{0}\theta_{0}nd_{0}}$ denoting range between the $n$-th antenna and the user. Moreover, $\theta_{0}=\cos\phi\in[-1,1]$ represents the spatial angle at the BS, with $\phi$ denoting the physical angle-of-departure (AoD) from the BS center to the user. Further, by means of Fresnel approximation, $r_{n}$ can be approximated as
@@ IV Received Beam Pattern of the Sparse DFT Codebook
In this section, we first introduce the sparse DFT codebook and then analyze its received beam pattern. Moreover, we design a periodical beam training codebook to reduce overhead and propose to activate a central sub-array for resolving the angular ambiguity. The main definitions in this section are given as follows, including the received beam pattern and angular support.
@@
$\Gamma_{\mu}^{\mathcal{L}}(r_{0},\theta_{0})=\theta_{\rm right}-\theta_{\rm left},$ (11)
@@ IV-A Sparse DFT Codebook
Each column of the DFT Codebook in (8) is a far-field channel response vector. For the sparse DFT codebook, we sample each column of the original DFT codebook using an interval of $U$, while the positions that are not sampled are padded with zeros. Then, $\bar{n}$-th column of the sparse DFT Codebook $\tilde{\mathbf{w}}_{\bar{n}}\in\mathbb{C}^{N\times 1}$ is given by $\tilde{\mathbf{w}}_{\bar{n}}=\bar{\mathbf{w}}_{\bar{n}}\odot\mathbf{n}(U)$, where
@@ V-B Phase 2: Angular Ambiguity Elimination
In the second phase, we propose an efficient method to resolve the angular ambiguity. The key idea is to utilize a central sub-array to sequentially examine the candidate angles in (26). Specifically, we activate a central sub-array with $M$ antennas to eliminate the angular ambiguity, for which the codebook ${{\mathcal{V}}}_{\rm{Sub}}$ is presented as follows
@@
It is noteworthy that there are two criteria for ensuring the effectiveness of angular ambiguity elimination. The first one is that the user should be located in the far-field region of the activated central sub-array to avoid the energy-spread effect. Moreover, considering that the interval between two adjacent candidate angles is $\frac{2}{U}$, extra angular ambiguity will be introduced if the beam width of the central sub-array is wider than $\frac{4}{U}$. Therefore, the second criterion is that the beam width of the central sub-array is required to be smaller than $\frac{4}{U}$, which ensures no interference between two adjacent candidate angles.
To satisfy the first criterion, we set the Rayleigh distance of the sub-array to be less than the Fresnel distance of the XL-array, which is given by
@@
For the second criterion, we have $\frac{4}{M}\leq\frac{4}{U}$, which can be simplified to $M\geq U$. Therefore, the number of antennas of the central sub-array needs to satisfy
@@
Moreover, from (30), the activation interval $U$ has the following constraint $U\leq\sqrt{1.2{(N-1)}}$. For example, given a setup where $N=257$ and $U=16$, we have $16\leq M\leq 17.5$. In Fig. 7, we plot the beam width of the sub-array with $M=17$. It is observed that two adjacent candidate user angles have no considerable mutual interference, which indicates the effectiveness of angular ambiguity elimination.
Figure 7: Illustration of angle ambiguity eliminated by the central sub-array, where $N=257$, $M=17$ and $f=30$ GHz. The beam is steered towards an candidate angle $1/16$. The red lines are two adjacent candidate user angles $-1/16$ and $1/16$.
Based on the above, the BS activates a central sub-array comprising $M$ antennas satisfying (30) to resolve the angular ambiguity. For the candidate user angles ${\theta}_{\check{s}}+u\frac{2}{U},\forall u\in\mathcal{U}$ in (26), the codewords chosen from (27) to resolve the angular ambiguity are $\mathbf{v}_{\check{s}+uQ},\forall u\in\mathcal{U}$. Then, the BS sequentially transmits $U$ pilot symbols with beamforming vectors $\{\mathbf{v}_{\check{s}+uQ}\}$ as illustrated in Fig. 5 (b). For each codeword, the received signal power at the user in the second phase is given by

### v2

@@ Abstract
Extremely large-scale arrays (XL-arrays) have emerged as one promising technology to improve the spectral efficiency and spatial resolution in future sixth generation (6G) wireless systems. The drastic increase in the number of antennas renders the communication users more likely to be located in the near-field region, which requires a more accurate spherical (instead of planar) wavefront propagation modeling. However, this also inevitably incurs unaffordable beam training overhead when performing a two-dimensional (2D) beam-search in both the angular and range domains. To address this issue, we first introduce in this paper a new sparse discrete Fourier transform (DFT) codebook, which exhibits the angular periodicity in the received beam pattern at the user. This thus motivates us to propose a three-phase beam training scheme. Specifically, in the first phase, we utilize the sparse DFT codebook for beam sweeping in an angular subspace and estimate candidate user angles according to the received beam pattern. Then, a central subarray is activated to scan specific candidate angles for resolving the issue of angular ambiguity for identifying the user angle. In the third phase, the polar-domain codebook is applied in the estimated angle to search the best effective user range. Finally, numerical results show that our proposed beam training scheme enabled by the sparse DFT codebook achieves 98.67% beam training overhead reduction as compared to the exhaustive-search scheme, yet without compromising rate performance in the high signal-to-ratio (SNR) regime.
@@ I Introduction
Extremely large-scale arrays/surfaces (XL-arrays/surfaces) have been envisioned as one of the key ingredients to drive the evolution of six generation (6G) wireless systems[1, 2, 3, 4]. Specifically, XL-arrays/surfaces with a significant number of antennas can be deployed at the base station (BS) to achieve ultra-high spectral efficiency and spatial resolution, hence accommodating the escalating demands for new applications such as spanning metaverse and digital twin [5, 6, 7]. The drastic increase in the number of antennas in high-frequency bands represents a qualitative paradigm shift in the electromagnetic (EM) propagation modeling, giving rise to the new near-field communications [8].
Particularly, different from the far-field EM propagation which is simply approximated by planar waves, the near-field channel modeling necessitates the use of more accurate spherical waves[9, 10, 11]. As such, near-field communications possess several unique properties in contrast to far-field communications. First, the spherical wavefront characteristic opens up the possibility of near-field beamfocusing, for which the beam energy can be concentrated at a specific location/region rather than a spatial angle typically for far-field beamforming[12, 13, 14]. The beam-focusing capability of near-field communications enables XL-array to flexibly form highly directional beams in both the angle and range domain, and hence can be leveraged in various applications to improve the system performance, such as mitigating the inter-user interference, improving the accuracy of sensing and localization[15, 16, 17], and enhancing the charging efficiency of wireless power transfer[18]. Second, the rank of line-of-sight (LoS) channels for near-field multiple-input multiple-output (MIMO) communication systems can be larger than one, hence enhancing the spatial multiplexing gains [19]. In this paper, we propose an efficient near-field beam training scheme with an sparse discrete Fourier transform (DFT) codebook by using the sparse antenna activation method, which significantly reduces the beam training overhead.
@@ I-A1 Near-Field Wireless Systems
Near-field communications bring new opportunities and challenges, which has motivated upsurging research interest. For example, in [12], the authors studied the beamfocusing design based on the fully-digital architectures, hybrid phase shifter-based precoders, and dynamic metasurface antenna architecture for XL-MIMO arrays. It is shown that near-field beamfocusing provides new degree of freedom (DoF) to mitigate the interference among users, even when they are located at the same spatial angle. In addition, a new concept of location division multiple access (LDMA) was proposed in [20], which exploits orthogonality of the near-field beamfocusing vectors in the range domain to serve different users at the same angle. The authors in [21] developed a framework for analyzing and designing XL-MIMO systems with spatial non-stationarity. It was revealed that the performance of the proposed framework approaches that of the conventional full-antenna array based designs albeit with lower complexity. In addition, a holographic metasurface antennas (HMAs) based multi-user system was investigated in [22], where the digital transmit precoder and the analog HMA weighting matrix were jointly optimized to minimize the transmit power. The authors in [23] considered a simultaneous wireless information and power transfer (SWIPT) system, where energy harvesting (EH) and information decoding receivers are located in the near- and far-field regions of the XL-array, respectively. The beam scheduling and power allocation were jointly optimized to maximize the weighted-sum power harvested at EH receivers [23]. Moreover, a directional modulation system was developed for near-field physical layer security systems[24]. Specifically, a fully analog precoding algorithm along with artificial noise and power allocation was proposed to realize secure transmission in both the angular and range domains. Near-field sensing with XL-array was studied in [25], where the closed-form expressions of the Cramér-Rao Bounds for both the angle and range estimations are derived.
@@ I-A2 Near-field Beam Training
In high-frequency bands, direct channel estimation methods may not be very effective due to severe path-loss and signal misalignment. As such, beam training is efficient in establishing initial links with high signal-to-noise ratio (SNR) for data transmission and channel state information (CSI) acquisition [26]. However, due to the spherical wave propagation, near-field beam training is more challenging compared with its far-field counterpart as it requires a joint beam search over both the angular and range domain. Particularly, conventional far-field beam training will suffer from significant performance loss in the near-field region due to the so-called energy-spread effect, for which the energy of a far-field beamformer is no longer steered towards one angle, but spread in multiple angles. Hence, the conventional far-field beam training method cannot be directly applied to the near-field beam training. To address this issue, the authors in [27] proposed a new codebook design in the polar domain, for which the angular domain is uniformly sampled whereas the range domain is non-uniformly sampled. One can simply invoke this codebook for the exhaustive-search based near-field beam training. However, the overhead of this scheme is the product of the number of antennas and range samples, which is prohibitively high for the implementation of XL-arrays. To reduce the overhead, the authors in [28] proposed a novel two-phase near-field beam training method. This method leveraged a key observation that the true user angle approximately lies in the middle of an angular support with high received SNRs. Based on this observation, one can first estimate the user angle by using far-field DFT beams, and then estimate the user range with the polar-domain codebook in [27]. Besides, deep learning techniques have also been exploited in [29] and [30] to reduce the near-field beam training overhead, where deep neural networks (DNN) are trained based on conventional far-field codebooks and near-field codebook, respectively. However, the training overhead of these methods scales linearly with the number of antennas, which is still unaffordable for communication systems. This issue motivates the design of efficient hierarchical beam training schemes for near-field communications to reduce the training overhead to the logarithmic order, e.g., [31, 32, 33]. However, hierarchical beam training schemes suffer from several inherent drawbacks. First, hierarchical methods usually require frequent feedback and neglect underlying transmission delays. Moreover, there exists the error propagation issue due to the progressive beam search [34].
@@ I-B Motivations and Contributions
The existing works on near-field beam training inevitably incur unacceptable overhead, while the off-grid channel estimation methods face highly computational complexity such as high dimensional matrix inversion. Moreover, the beam training overhead of these existing works scales linearly with the number of antennas apart from hierarchical beam training schemes, which suffers from several inherent drawbacks such as error propagation and user feed-back overhead [34]. Motivated by the above, this paper explores a new sparse DFT codebook and a three-phase beam training scheme to reduce the near-field beam training overhead, which scales with the square root of the number of antennas. The main contributions are summarized as follows.
  • First, we propose a novel sparse DFT codebook by sparsely activating the XL-array antennas and constructing the reduced DFT codebook with the equivalent sparse linear array (SLA). Specifically, the sparse DFT codebook consists of sparse far-field channel response vectors, which is sampled from far-field channel response vectors. Then, we characterize the received beam pattern at the near-field user when the sparse DFT codebook is used for beam sweeping. Interestingly, it is shown that the received beam pattern exhibits periodicity in the angular domain, while there still exists the energy-spread effect. Then, we show that the user angle information is contained in a period of the received beam pattern at the user and can be estimated via a defined angular support.
  • Second, we propose a novel three-phase beam training scheme based on the sparse DFT codebook. Specifically, in the first phase, we utilize a small number of the sparse DFT codewords to sweep an angular subspace and estimate one candidate user angle according to the middle of the defined angular support. Then, in the second phase, we activate a central subarray to resolve the angular ambiguity by virtue of the periodicity of the received beam pattern at the user. Subsequently, the polar-domain codebook is utilized to search the best user range in the estimated user angle.
  • Finally, extensive numerical results are presented to demonstrate the effectiveness of our proposed beam training scheme enabled by the sparse DFT codebook. It is shown that the proposed three-phase beam training scheme can achieve nearly the same performance with the exhaustive-search beam training method in the high-SNR regime, while reducing more than 98% of the training overhead. In the low-SNR regime, the proposed scheme suffers from slight performance loss, while the effective rate still significantly exceeds all benchmark schemes due to lower beam training overhead.
@@ I-C Organization and Notations
$D$ Array aperture size $M$ Number of antennas of the activated subarray
@@
${{\mathcal{V}}}_{\rm{Sub}}$ DFT codebook for the activated subarray $\bar{\mathcal{X}}_{\rm{Pol}}$ Polar-domain codebook
@@ II System Model
We consider a single-user XL-array downlink communication system, where the BS is equipped with a dense uniform linear array (ULA) with $N$ antennas. In this section, the near-field channel and signal model for the ULA are introduced.
@@ II-A Near-field Channel Model
which includes one LoS path $\mathbf{h}_{\rm LoS}^{H}$ and $L$ non-LoS (NLoS) paths. Herein, the parameters $r_{0}$ (${\bar{r}_{\ell}}$) and $\theta_{0}$ ($\bar{\theta}_{\ell}$) represent the range and spatial angle of the LoS ($\ell$-th NLoS) signal path. Moreover, the parameters $\beta$ and $\beta_{\ell}$ denote the LoS path and $\ell$-th NLoS path gain, respectively. Mathematically, $\beta$ can be modeled as [37]
$\beta=\sqrt{\frac{\kappa}{\kappa+1}}\frac{\sqrt{\beta_{0}}}{r_{{0}}}e^{-\frac{\jmath 2\pi r_{{0}}}{\lambda}},$ (2)
where $\kappa$ and $\beta_{0}$ represent the Rician factor and reference channel gain at a range of 1 m, respectively.
@@
with $r_{n}=\sqrt{r_{0}^{2}+{n}^{2}d_{0}^{2}-2r_{0}\theta_{0}nd_{0}}$ denoting range between the $n$-th antenna and the user. Moreover, $\theta_{0}=\cos\phi_{0}\in[-1,1]$ represents the spatial angle at the BS, with $\phi_{0}$ denoting the physical angle-of-departure (AoD) from the BS center to the user. Further, by means of Fresnel approximation, $r_{n}$ can be approximated as
@@ IV Received Beam Pattern of the Sparse DFT Codebook
In this section, we first introduce the sparse DFT codebook and then analyze its received beam pattern. Moreover, we design a periodical beam training codebook to reduce overhead and propose to activate a central subarray for resolving the angular ambiguity. The main definitions in this section are given as follows, including the received beam pattern and angular support.
@@
$\Gamma_{\mu}^{\mathcal{L}}(r_{0},\theta_{0})=\theta_{\rm right}-\theta_{\rm left}.$ (12)
@@ IV-A Sparse DFT Codebook
Each column of the DFT Codebook in (9) is a far-field channel response vector. For the sparse DFT codebook, we sample each column of the original DFT codebook using an interval of $U$, while the positions that are not sampled are padded with zeros. As such, $\bar{n}$-th column of the sparse DFT Codebook $\tilde{\mathbf{w}}_{\bar{n}}\in\mathbb{C}^{N\times 1}$ is given by $\tilde{\mathbf{w}}_{\bar{n}}=\bar{\mathbf{w}}_{\bar{n}}\odot\mathbf{n}(U)$, where
@@ V-B Phase 2: Angular Ambiguity Elimination
In the second phase, we propose an efficient method to resolve the angular ambiguity. The key idea is to utilize a central subarray to sequentially examine the candidate angles in (27). Specifically, we activate a central subarray with $M$ antennas to eliminate the angular ambiguity, for which the codebook ${{\mathcal{V}}}_{\rm{Sub}}$ is presented as follows
@@
It is noteworthy that there are two criteria for ensuring the effectiveness of angular ambiguity elimination. The first one is that the user should be located in the far-field region of the activated central subarray to avoid the energy-spread effect. Moreover, considering that the interval between two adjacent candidate angles is $\frac{2}{U}$, extra angular ambiguity will be introduced if the beam width of the central subarray is wider than $\frac{4}{U}$. Therefore, the second criterion is that the beam width of the central subarray is required to be smaller than $\frac{4}{U}$, which ensures no interference between two adjacent candidate angles.
To satisfy the first criterion, we set the Rayleigh distance of the subarray to be less than the Fresnel distance of the XL-array, which is given by
@@
For the second criterion, we have
$\frac{4}{M}\leq\frac{4}{U},$ (31)
which can be simplified to $M\geq U$. Therefore, the number of antennas of the central subarray needs to satisfy
@@
Moreover, from (32), the activation interval $U$ has the following constraint $U\leq\sqrt{1.2{(N-1)}}$. For example, given a setup where $N=257$ and $U=16$, we have $16\leq M\leq 17.5$. In Fig. 7, we plot the beam width of the subarray with $M=17$. It is observed that two adjacent candidate user angles have no considerable mutual interference, which indicates the effectiveness of angular ambiguity elimination.
Figure 7: Illustration of angle ambiguity eliminated by the central subarray, where $N=257$, $M=17$ and $f=30$ GHz. The beam is steered towards an candidate angle $1/16$. The red lines are two adjacent candidate user angles $-1/16$ and $1/16$.
Based on the above, the BS activates a central subarray comprising $M$ antennas satisfying (32) to resolve the angular ambiguity. For the candidate user angles ${\theta}_{\check{s}}+u\frac{2}{U},\forall u\in\mathcal{U}$ in (27), the codewords chosen from (28) to resolve the angular ambiguity are $\mathbf{v}_{\check{s}+uQ},\forall u\in\mathcal{U}$. Then, the BS sequentially transmits $U$ pilot symbols with beamforming vectors $\{\mathbf{v}_{\check{s}+uQ}\}$ as illustrated in Fig. 5 (b). For each codeword, the received signal power at the user in the second phase is given by